OpenStudy (anonymous):

Simplify the following expression. sqrt -49 sqrt -12 *** there is no comma seperating the 2 square roots

OpenStudy (anonymous):

\[\sqrt{-49}=7i\] is a start

OpenStudy (anonymous):

\[\sqrt{-12}=2\sqrt{3}i\]

OpenStudy (anonymous):

multiply together and recall that \(i^2=-1\)

OpenStudy (anonymous):

so what would you get as the final outcome?